# Rigidity results for initial data sets satisfying the dominant energy condition

### v3

@@ 5. Approximating a compact, convex polytope by smooth domains
$\leq(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0$
as $l\to\infty$. On the other hand, it follows from Proposition 5.7 that
$\sup_{p\in\mathbb{R}^{n}}\sup_{(\delta_{l}\lambda_{l})^{-1}\leq r\leq 1}\bigg{(}r^{\sigma+1-n}\int_{\Sigma_{\lambda_{l}}\cap B_{r}(p)}\max\{-W_{\lambda_{l}},0\}^{\sigma}\bigg{)}^{\frac{1}{\sigma}}\to 0$
as $l\to\infty$. Putting these facts together, the assertion follows. ∎
@@ 6. Proof of Theorem II
Let $\Omega$, $g$, $q$, $H$ satisfy the assumptions in Theorem II. Furthermore assume that $\Omega$ satisfies the Matching Angle hypothesis. We obtain the spinor bundle $\mathcal{S}$ and the twisted spinor bundle $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ over an open set containing $\Omega$.
Let $U$ denote a Euclidean ball such that the closure of $U$ is contained in the interior of $\Omega$. Consider a sequence $\lambda_{l}\to\infty$. Given $l$, let $N^{(l)}\colon\Sigma_{\lambda_{l}}\to S^{n-1}$ denote the map from (5.1). Recall that $N^{(l)}$ has degree $1$. Note that $U\subset\Omega_{\lambda_{l}}$ if $l$ is sufficiently large.
@@
Corollary 6.2.
where $C$ is a constant that does not depend on $l$.
Proof.
If we choose $K$ to be a large multiple of $\sup_{\Omega}|q|$, then we can bound
$(|\nabla^{\mathcal{E}}s|-K\,|s|)_{+}\leq|\nabla^{\mathcal{E},q}s|.$
Putting these facts together, the assertion follows. ∎
@@
Proposition 6.5.
We have
for every smooth function $F$ on $\Omega_{\lambda_{l}}$.
Proof.
This follows from Corollary 5.8 together with the Fefferman-Phong estimate (cf. [Brendle], Appendix A). ∎
@@
The following assertions hold.
@@
    $q^{\rm tan}=0$ along each of the boundary faces.
Having established Proposition 6.7, the arguments in Subsection 4.3 imply that $(\Omega,g,q)$ can be isometrically immersed as a spacelike hypersurface in the Minkowski spacetime $\mathbb{R}^{n,1}$ such that $q$ is the second fundamental form.
@@
Suppose that $n\geq 3$ is an odd integer. By Proposition 3.8, we can find a non-trivial section $s^{(l)}\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
@@
if $l$ is sufficiently large. Using Corollary 6.2, we obtain
@@
By scaling, we can arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
@@
After passing to a subsequence if necessary, the sequence $s^{(l)}$ converges weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$. The limit $s$ is defined on the interior of $\Omega$ and satisfies $\int_{U}|s|^{2}=1$. Using (6.1), we obtain $\nabla^{\mathcal{E},q}s=0$. In particular, $s$ is smooth. Using $\nabla^{\mathcal{E},q}$-parallel transport along radial lines emanating from some point in the interior of $\Omega\subset\mathbb{R}^{n}$, we extend $s$ to a smooth section which is defined on an open set containing $\Omega$ and satisfies $\nabla^{\mathcal{E},q}s=0$ at each point in $\Omega$.
@@
$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0$
as $l\to\infty$.
@@
Since $\nabla^{\mathcal{E},q}s=0$, we conclude that
@@
The assertion now follows from (6.1) and the fact that $s^{(l)}\to s$ weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$. ∎
@@
$\int_{\Sigma_{\lambda_{l}}}|\chi^{(l)}s-s|^{2}\to 0$
as $l\to\infty$. Here, $\chi^{(l)}$ denotes the boundary operator on $\Sigma_{\lambda_{l}}$.
@@
Recall that $s^{(l)}$ satisfies the boundary condition $\chi^{(l)}s^{(l)}=s^{(l)}$ at each point on $\Sigma_{\lambda_{l}}$. Hence, the assertion follows from Lemma 6.8. ∎
For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is non-empty. Moreover, it follows from Lemma 6.9 that $s$ satisfies the boundary condition
$\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(N_{i_{0}})$
on $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$.
@@
As above, we define
$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ each point on $\Omega$}\}.$
@@
This follows from Lemma 6.11 together with the fact that $\text{\rm span}\{N_{i}:i\in I\}=\mathbb{R}^{n}$. ∎
@@
on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$. In the next step, we differentiate this relation in tangential direction. Arguing as in the proof of Lemma 4.2, we obtain
$\gamma(h(X))\circ s+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\circ s=0$
whenever $p$ is a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ and $X$ is a tangent vector to that boundary face. Using Lemma 6.12, we conclude that
$\gamma(h(X))+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$
whenever $p$ is a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ and $X$ is a tangent vector to that boundary face. Note that $i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\in\text{\rm End}(\mathcal{S}|_{p})$ is self-adjoint, while $\gamma(h(X))\in\text{\rm End}(\mathcal{S}|_{p})$ is skew-adjoint. Thus, we conclude that
$\gamma(h(X))=0\in\text{\rm End}(\mathcal{S}|_{p}),\quad\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$
whenever $p$ is a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ and $X$ is a tangent vector to that boundary face. From this, the assertion follows. ∎
@@
Suppose now that $n\geq 2$ is an even integer. By Corollary 3.12, we can find a non-trivial section $s^{(l)}\in H^{1}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
@@
Moreover, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}^{(l)})$ on $\Sigma_{\lambda_{l}}$, where the map $\hat{N}^{(l)}:\Sigma_{\lambda_{l}}\to S^{n}$ is defined by $\hat{N}^{(l)}(x)=\varphi(N^{(l)}(x),\hat{t}_{l})$ for all $x\in\Sigma_{\lambda_{l}}$. By assumption, $\mu-|J|\geq 0$. Using Corollary 6.6, we obtain
@@
if $l$ is sufficiently large. Using Corollary 6.2, we obtain
@@
By scaling, we can arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
@@
and
$\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C.$
After passing to a subsequence if necessary, we may assume that $\hat{t}_{l}\to\hat{t}$. Moreover, the sequence $s^{(l)}$ converges weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$. The limit $s$ is defined on the interior of $\Omega$ and satisfies $\int_{U}|s|^{2}=1$. Using (6.2), we obtain $\nabla^{\mathcal{E},q}s=0$. In particular, $s$ is smooth. Using $\nabla^{\mathcal{E},q}$-parallel transport along radial lines emanating from some point in the interior of $\Omega\subset\mathbb{R}^{n}$, we extend $s$ to a smooth section which is defined on an open set containing $\Omega$ and satisfies $\nabla^{\mathcal{E},q}s=0$ at each point in $\Omega$.
@@
$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0$
as $l\to\infty$.
@@
$\int_{\Sigma_{\lambda_{l}}}|\gamma(\nu)\circ s-s\circ\gamma_{0}(\hat{N}^{(l)})|^{2}\to 0$
as $l\to\infty$. Here, $\chi^{(l)}$ denotes the boundary operator on $\Sigma_{\lambda_{l}}$.
@@
Recall that $s^{(l)}$ satisfies the boundary condition $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}^{(l)})$ on $\Sigma_{\lambda_{l}}$ at each point on $\Sigma_{\lambda_{l}}$. Hence, the assertion follows from Lemma 6.14. ∎
For each $i\in I$, we define $\hat{N}_{i}=\varphi(N_{i},\hat{t})\in S^{n}$. For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is non-empty. Moreover, it follows from Lemma 6.15 that $s$ satisfies the boundary condition
@@
on $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$. By Lemma 3.2 in [Brendle], the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is a dense subset of $\Omega\cap\{u_{i_{0}}=0\}$. Thus, we conclude that
@@
We consider a point $p\in\Omega\cap\{u_{i_{1}}=0\}\cap\{u_{i_{2}}=0\}$. Then
@@
at the point $p$. On the other hand, the Matching Angle Hypothesis implies that $\langle\nu_{i_{1}},\nu_{i_{2}}\rangle_{g}\,|s|^{2}=\langle N_{i_{1}},N_{i_{2}}\rangle\,|s|^{2}$ at the point $p$. Putting these facts together, we conclude that $\langle N_{i_{1}},N_{i_{2}}\rangle\,|s|^{2}=\langle\hat{N}_{i_{1}},\hat{N}_{i_{2}}\rangle\,|s|^{2}$ at the point $p$. Since $|s|^{2}>0$ at each point in $\Omega$, the assertion follows. ∎
@@
Suppose that the assertion is false. By Lemma 3.9, the map $\varphi(\cdot,\hat{t})$ is Lipschitz continuous with a Lipschitz constant strictly less than $1$. This contradicts Lemma 6.16. ∎
@@
We again define
$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ each point on $\Omega$}\}.$
@@
Note that $\text{\rm span}\{N_{i}:i\in I\}=\mathbb{R}^{n}$. Using Lemma 6.17, it follows that $\text{\rm span}\{\hat{N}_{i}:i\in I\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}$. By Lemma 6.19, $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n}$. Since the volume element acts as a scalar, we deduce that $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n+1}$. Arguing as in the proof of Lemma 4.6, the assertion follows. ∎

### v4

@@ 5. Approximating a compact convex polytope by smooth domains
$\leq(\delta_{l}\lambda_{l})^{-1}\sup_{\Sigma_{\lambda_{l}}}\max\{-W_{\lambda_{l}},0\}\to 0\quad\text{ as }l\to\infty.$
On the other hand, it follows from Proposition 5.8 that
$\sup_{p\in\mathbb{R}^{n}}\sup_{(\delta_{l}\lambda_{l})^{-1}\leq r\leq 1}\bigg(r^{\sigma+1-n}\int_{\Sigma_{\lambda_{l}}\cap B_{r}(p)}\max\{-W_{\lambda_{l}},0\}^{\sigma}\bigg)^{\frac{1}{\sigma}}\to 0\quad\text{ as }l\to\infty.$
Putting these facts together, the assertion follows. ∎
@@ 6. Proof of Theorem II
Throughout this section, we assume that $\Omega$, $g$, $q$ satisfy the assumptions in Theorem II. In particular, we assume that $\Omega$ satisfies the Matching Angle hypothesis. We consider the spinor bundle $\mathcal{S}$ and the twisted spinor bundle $\mathcal{E}=\mathcal{S}\otimes S_{0}^{*}$ over an open set containing $\Omega$.
Let $U$ denote a Euclidean ball such that the closure of $U$ is contained in the interior of $\Omega$. Consider a sequence $\lambda_{l}\to\infty$. Given $l$, let $N^{(l)}\colon\Sigma_{\lambda_{l}}\to S^{n-1}$ denote the map defined in (5.1). Recall that $N^{(l)}$ has degree $1$. Note that $U\subset\Omega_{\lambda_{l}}$ if $l$ is sufficiently large.
@@
Proposition 6.3.
We have
for every smooth function $F$ on $\Omega_{\lambda_{l}}$.
Proof.
The statement follows by combining [Brendle]*Corollary A.7 (see also [Fefferman-Phong]) with Corollary 5.9. ∎
Corollary 6.4.
where $C$ is a constant that does not depend on $l$.
Proof.
By the triangle inequality, we can bound $|\nabla^{\mathcal{E}}s|\leq|\nabla^{\mathcal{E},q}s|+K\,|s|$, where $K$ is a constant that does not depend on $l$. This implies $(|\nabla^{\mathcal{E}}s|-K\,|s|)_{+}\leq|\nabla^{\mathcal{E},q}s|$.
Putting these facts together, the assertion follows. ∎
@@
The following statements hold.
@@
    Along each of the boundary faces, we have $q^{\rm tan}=0$.
Having established Proposition 6.7, the arguments in Section 4.3 imply that $(\Omega,g,q)$ can be isometrically immersed as a spacelike hypersurface in the Minkowski spacetime $\mathbb{R}^{n,1}$, with second fundamental form equal to $q$.
@@
Suppose that $n\geq 3$ is an odd integer. For each $l$, Proposition 3.7 implies the existence of a non-trivial section $s^{(l)}\in C^{\infty}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
@@
if $l$ is sufficiently large. Using this estimate together with Corollary 6.4, we obtain
@@
By scaling, we may arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
@@
Passing to a subsequence if necessary, the sequence $s^{(l)}$ converges weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$. The limit $s$ is defined on the interior of $\Omega$ and satisfies $\int_{U}|s|^{2}=1$. Using (6.1), we obtain $\nabla^{\mathcal{E},q}s=0$. In particular, $s$ is smooth. Using $\nabla^{\mathcal{E},q}$-parallel transport along radial lines emanating from some point in the interior of $\Omega\subset\mathbb{R}^{n}$, we extend $s$ to a smooth section which is defined on an open set containing $\Omega$ and satisfies $\nabla^{\mathcal{E},q}s=0$ at each point in $\Omega$.
@@
$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0\quad\text{ as }l\to\infty.$
@@
Since $\nabla^{\mathcal{E},q}s=0$, it follows that
@@
The assertion now follows from (6.1) together with the fact that $s^{(l)}\to s$ in $L_{\text{\rm loc}}^{2}(\Omega\setminus\partial\Omega)$. ∎
@@
$\int_{\Sigma_{\lambda_{l}}}|\chi^{(l)}s-s|^{2}\to 0\quad\text{ as }l\to\infty.$
Here, $\chi^{(l)}$ denotes the boundary operator on $\Sigma_{\lambda_{l}}$.
@@
Recall that $s^{(l)}$ satisfies the boundary condition $\chi^{(l)}s^{(l)}=s^{(l)}$ at each point on $\Sigma_{\lambda_{l}}$. Hence, the assertion follows from Proposition 6.8. ∎
For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is nonempty (see [Brendle]*Lemma 3.2). Moreover, it follows from Corollary 6.9 that $s$ satisfies the boundary condition $\gamma(\nu_{i_{0}})\circ s=s\circ\gamma_{0}(N_{i_{0}})$ on $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$.
@@
As above, we define $L:=\{\sigma\in S_{0}\colon\text{\rm$s\sigma=0$ at each point in $\Omega$}\}$.
@@
This follows from Lemma 6.11 together with the fact that
$\text{\rm span}\{N_{i}\colon i\in I\}=\mathbb{R}^{n}.\qed$
@@
on the boundary face
$\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}.$
In the next step, we differentiate this identity in tangential direction. Let $p$ be a point on the boundary face $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$, and let $X$ be a tangent vector to that boundary face. Arguing as in the proof of Lemma 4.2, we obtain
$\gamma(h(X))\circ s+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\circ s=0.$
Since $s\in\text{\rm Hom}(S_{0},\mathcal{S}|_{p})$ is invertible by Lemma 6.12, it follows that
$\gamma(h(X))+i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p}).$
Note that $i\,\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)\in\text{\rm End}(\mathcal{S}|_{p})$ is self-adjoint, while $\gamma(h(X))\in\text{\rm End}(\mathcal{S}|_{p})$ is skew-adjoint. Thus, we conclude that
$\gamma(h(X))=0\in\text{\rm End}(\mathcal{S}|_{p})\quad\text{ and }\quad\gamma(q(X)^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p}).$
From this, the assertion follows. ∎
@@
Suppose that $n\geq 2$ is an even integer. For each $l$, Corollary 3.11 implies the existence of a non-trivial section $s^{(l)}\in H^{1}(\Omega_{\lambda_{l}},\mathcal{E})$ such that
@@
$\quad+\tfrac{1}{2}\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}.$
Moreover, for each $l$, we can find an element $\hat{t}_{l}\in S^{1}$ such that $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}^{(l)})$ on $\Sigma_{\lambda_{l}}$, where the map $\hat{N}^{(l)}\colon\Sigma_{\lambda_{l}}\to S^{n}$ is defined by $\hat{N}^{(l)}(x)=\varphi(N^{(l)}(x),\hat{t}_{l})$ for all $x\in\Sigma_{\lambda_{l}}$. By assumption, $\mu-|J|\geq 0$. Using Corollary 6.6, we obtain
$\int_{\Sigma_{\lambda_{l}}}(|\mathrm{tr}(q)-q(\nu,\nu)|$ $+\|dN^{(l)}\|_{\mathrm{tr}}-H)\,|s^{(l)}|^{2}\leq o(1)\int_{\Omega_{\lambda_{l}}}|\nabla^{\mathcal{E},q}s^{(l)}|^{2}+o(1)\int_{U}|s^{(l)}|^{2}.$
@@
if $l$ is sufficiently large. Using this estimate together with Corollary 6.4, we obtain
@@
By scaling, we may arrange that $\int_{U}|s^{(l)}|^{2}=1$ if $l$ is sufficiently large. Then
@@
and $\int_{\Omega_{\lambda_{l}}}|s^{(l)}|^{2}\leq C$. After passing to a subsequence if necessary, we may assume that $\hat{t}_{l}\to\hat{t}$ for some $\hat{t}\in S^{1}$. Moreover, the sequence $s^{(l)}$ converges weakly in $H_{\text{\rm loc}}^{1}(\Omega\setminus\partial\Omega)$. The limit $s$ is defined on the interior of $\Omega$ and satisfies $\int_{U}|s|^{2}=1$. Using (6.2), we obtain $\nabla^{\mathcal{E},q}s=0$. In particular, $s$ is smooth. Using $\nabla^{\mathcal{E},q}$-parallel transport along radial lines emanating from some point in the interior of $\Omega\subset\mathbb{R}^{n}$, we extend $s$ to a smooth section which is defined on an open set containing $\Omega$ and satisfies $\nabla^{\mathcal{E},q}s=0$ at each point in $\Omega$.
@@
$\int_{\Sigma_{\lambda_{l}}}|s^{(l)}-s|^{2}\to 0\quad\text{ as }l\to\infty.$
@@
$\int_{\Sigma_{\lambda_{l}}}|\gamma(\nu)\circ s-s\circ\gamma_{0}(\hat{N}^{(l)})|^{2}\to 0\quad\text{ as }l\to\infty.$
@@
Recall that $s^{(l)}$ satisfies the boundary condition $\gamma(\nu)\circ s^{(l)}=s^{(l)}\circ\gamma_{0}(\hat{N}^{(l)})$ on $\Sigma_{\lambda_{l}}$ at each point on $\Sigma_{\lambda_{l}}$. Hence, the assertion follows from Proposition 6.14. ∎
For each $i\in I$, we define $\hat{N}_{i}=\varphi(N_{i},\hat{t})\in S^{n}$. For each $i_{0}\in I$, Assumption 5.1 implies that the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is nonempty (see [Brendle]*Lemma 3.2). Moreover, it follows from Corollary 6.15 that $s$ satisfies the boundary condition
@@
on $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$. By [Brendle]*Lemma 3.2, the set $\{u_{i_{0}}=0\}\cap\bigcap_{i\in I\setminus\{i_{0}\}}\{u_{i}<0\}$ is a dense subset of $\Omega\cap\{u_{i_{0}}=0\}$. Thus, we conclude that
@@
By assumption, we can find a point $p\in\Omega\cap\{u_{i_{1}}=0\}\cap\{u_{i_{2}}=0\}$. Then
@@
at the point $p$. On the other hand, the Matching Angle Hypothesis implies that
$\langle\nu_{i_{1}},\nu_{i_{2}}\rangle_{g}\,|s|^{2}=\langle N_{i_{1}},N_{i_{2}}\rangle\,|s|^{2}$
at the point $p$. Putting these facts together, we conclude that $\langle N_{i_{1}},N_{i_{2}}\rangle\,|s|^{2}=\langle\hat{N}_{i_{1}},\hat{N}_{i_{2}}\rangle\,|s|^{2}$ at the point $p$. Since $|s|^{2}>0$ at each point in $\Omega$, the assertion follows. ∎
@@
[…]
$|N_{i_{1}}-N_{i_{2}}|^{2}>|\varphi(N_{i_{1}},\hat{t})-\varphi(N_{i_{2}},\hat{t})|^{2}=|\hat{N}_{i_{1}}-\hat{N}_{i_{2}}|^{2}$
for $i_{1}\neq i_{2}$. This contradicts Lemma 6.16. ∎
@@
We again define $L:=\{\sigma\in S_{0}\colon\text{\rm$s\sigma=0$ at each point in $\Omega$}\}$.
@@
Note that $\text{\rm span}\{N_{i}\colon i\in I\}=\mathbb{R}^{n}$. Using Lemma 6.17, it follows that
$\text{\rm span}\{\hat{N}_{i}\colon i\in I\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}.$
By Lemma 6.19, $L$ is invariant under $\gamma_{0}(\xi)$ for all $\xi\in\mathbb{R}^{n}$. Since the volume element acts as a scalar multiple of the identity, we deduce that $L$ is invariant under $\gamma_{0}(\xi)$ for all $\xi\in\mathbb{R}^{n+1}$. Arguing as in the proof of Lemma 4.5, the assertion follows. ∎